# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 1. Introduction
Random graph models based on preferential attachment are quite well-understood nowadays due to a large body of research taken place over the last three decades (see [36, 37] and the references therein for a good overview of the literature). These models have become a staple in random graph theory due to their use in modelling real-world networks. Still, a clear gap between most of these models and networks that form in the real world is the fact that preferential attachment graphs allow for growth only. Vertices and edges are sequentially attached to the graph, causing the graph to grow. However, this is not particularly realistic, as nodes and bonds in real-world networks can be removed in almost all contexts as well. As an example, people are born but also pass away; friendships are formed but can also be broken; users of social media can befriend or follow others, but also unfriend or unfollow them and even remove their accounts all together.
@@
Some recent work partially addresses these concerns. Diaz, Lichev, and the author study a model with uniform attachment and uniform vertex removal ($m$ is random, $p_{2}=p_{4}=0$), where the local weak limit, existence of a giant component, and the size and location of the maximum degree are studied [17]. Work of Bellin and Kammerer et al. [5, 6, 27] studies a tree model of uniform attachment and uniform vertex removal ($m=1$, $p_{2}=p_{4}=0$), but where vertices are not removed but are ‘frozen’ and can then no longer make new connections, and where the choice to freeze vertices need not be random. This allows them to study the model in a ‘critical window’ where they obtain precise results for the height of the tree and its scaling limit.
@@
The notion of persistence in evolving random graphs has been studied from different perspectives, starting with the work of Dereich and Mörters [16]. Later, conditions under which persistence and the lack thereof hold have been weakened or omitted by Galashin [20], Banerjee and Bhamidi [2], and Iyer [25]. In short (and omitting minor technical assumptions), using the PAVD formulation with $d\equiv 0$, so that there is no vertex death and we recover the classic preferential attachment model, persistence of the maximum degree occurs almost surely if and only if
@@
Heuristically, persistence occurs when high-degree vertices have a sufficiently strong advantage over low-degree vertices to establish new connections, so they are most likely to increase their in-degree further. This allows one vertex to attain the largest degree for all but finitely many steps, instead of the largest degree switching between vertices infinitely often. The transition between whether the advantage is strong enough or not lies exactly at the point when the series in (1.3) goes from finite to infinity.
@@
Our contribution. In this article we study the two regimes outlined above. In the infinite lifetime regime, in which vertices can live forever with positive probability, we provide conditions, similar to (1.3), under which a persistent $m$-hub exists almost surely conditionally on the tree process $(T_{n})_{n\in\mathbb{N}}$ surviving, for each $m\in\mathbb{N}$, as well as conditions under which persistent hubs do not exist. This analysis generalises and expands recent work of Iyer [25], which focusses on the case $d\equiv 0$ and $m=1$ and provides a novel perspective used to analyse the (non-)existence of persistent hubs.
@@
The second regime is the finite lifetime regime, in which vertices are killed after a finite number of steps almost surely. Here, we provide conditions, similar to (1.3), under which persistence occurs. This builds on recent work by the author and Heydenreich [23], in which lack of persistence is studied for the PAVD model in both the infinite and the finite lifetime regime, and extends the methodology developed by Deijfen in [13] using an embedding of the discrete tree process in a continuous-time branching process known as a Crump-Mode-Jagers branching process. We use a more precise formulation of this embedding compared to Deijfen, which allows us to simplify certain proofs of results in [13], but mainly to develop novel and precise results regarding the behaviour of the branching process embedding that were unattainable previously. Furthermore, we extend ideas used by Banerjee and Bhamidi in [2] to precisely analyse the optimal window in which the vertex with the largest degrees are found in the finite lifetime regime.
@@
Structure of the paper. We state the main results and the necessary assumptions in Section 2. The methodology used in the analysis in presented in Section 3. Section 4 studies the existence of persistent $m$-hubs in the infinite lifetime regime, whereas Section 5 focusses on the occurrence of persistence in the finite lifetime regime. We conclude with a discussion of our work, state several problems, and discuss possibilities for future directions in Section 6.
@@
Notation. Throughout the paper we use the following notation: we let :=$\mathbb{N}:=\{1,2,\ldots\}$ denote the natural numbers, set $\mathbb{N}_{0}:=\{0,1,\ldots\}$ to include zero and let $[t]:=\{i\in\mathbb{N}:i\leq t\}$ for any $t\geq 1$. For $x,y\in\mathbb{R}$, we let $\lceil x\rceil:=\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor:=\sup\{n\in\mathbb{Z}:n\leq x\}$, and let $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$. For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}$ such that $b_{n}$ is positive for all $n$ we say that $a_{n}=o(b_{n})$ and $a_{n}=\mathcal{O}(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$ and if there exists a constant $C>0$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$, respectively. We write $a_{n}=\Theta(b_{n})$ if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$. For random variables $X,(X_{n})_{n\in\mathbb{N}}$, and $Y$ we let $X_{n}\overset{\mathrm{d}}{\longrightarrow}X,X_{n}\overset{\mathbb{P}}{\longrightarrow}X$ and $X_{n}\overset{\mathrm{a.s.}}{\longrightarrow}X$ denote convergence in distribution, probability and almost sure convergence of $X_{n}$ to $X$, respectively. $X\preceq Y$ (resp. $X\succeq Y$) denotes that $X$ is stochastically dominated by $Y$ (resp. $X$ stochastically dominates $Y$). For a sequence $(\mathcal{E}_{t})_{t\in\mathbb{I}}$ of events, where $\mathbb{I}=\mathbb{N}$ or $\mathbb{I}=[0,\infty)$, we say that $\mathcal{E}_{t}$ holds with high probability when $\lim_{t\to\infty}\mathbb{P}\!\left(\mathcal{E}_{t}\right)=1$. Finally, for an event $\mathcal{S}$ such that $\mathbb{P}\!\left(\mathcal{S}\right)>0$ we let $\mathbb{P}_{\mathcal{S}}\left(\cdot\right):=\mathbb{P}(\cdot\,|\,\mathcal{S})$ and $\mathbb{E}_{\mathcal{S}}{}{\cdot}:=\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ denote the conditional probability measure and conditional expectation, respectively. Then, we let $\xrightarrow{\mathbb{P}_{\mathcal{S}}-\mathrm{a.s.}}$ respectively $\overset{\mathbb{P}_{\mathcal{S}}}{\longrightarrow}$ denote almost sure convergence and convergence in probability with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$.
@@ 2. Results
$\mathcal{S}:=\bigcap_{n\in\mathbb{N}}\{\mathcal{A}_{n}\neq\varnothing\}$ (2.1)
@@
$I_{n}^{(1)}:=\min\{i\in\mathcal{A}_{n}:\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\},$ (2.3)
@@
$I_{n}^{(m)}:=\min\{i\in\mathcal{A}_{n}\setminus\{I_{n}^{(1)},\ldots,I_{n}^{(m-1)}\}:\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\setminus\{I_{n}^{(1)},\ldots,I_{n}^{(m-1)}\}\}.$ (2.4)
@@
We are interested in the long-term behaviour of $I_{n}^{(m)}$ and $O_{n}$. First, we say that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent elder $\mathbb{P}_{\mathcal{S}}$-almost surely when there exists a $\mathbb{P}_{\mathcal{S}}$-almost surely finite random variable $O$ such that
@@
We say that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, when there exists a random variable $I^{(m)}$ such that
[…]
We say that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
$\Big{(}\frac{I_{n}^{(1)}}{O_{n}}\Big{)}_{n\in\mathbb{N}}\text{ is a tight sequence of random variables with respect to $\mathbb{P}_{\mathcal{S}}$},$ (P)
@@
Clearly, when the $1$-hub persists as in (PH-$m$) with $m=1$, then the existence of a persistent elder as in (PE) and persistence as in (P) are implied, since $1\leq O_{n}\leq I_{n}^{(1)}$ and $O_{n}$ is non-decreasing. When $m$-hubs do not persist for any $m\in\mathbb{N}$, the vertex that attains the $m^{\mathrm{th}}$ largest degree changes infinitely often. Persistence (and lack thereof) quantifies the rate of change of the maximum-degree vertex. Persistence tells us that the rate of change of $I_{n}^{(1)}$ is of the same rate of $O_{n}$, so that the maximum-degree vertex has a label (or ‘age’) of the same order as $O_{n}$, whereas lack of persistence tells us that the rate of change of $I_{n}^{(1)}$ is faster than that of $O_{n}$.
@@
  • A class of PAVD models for which we prove that it does not contain an $m$-persistent hub for any $m\in\mathbb{N}$, but persistence does occur. (Theorem 2.6)
This complements recent work of the author and Heydenreich in [23], where a class of PAVD models is studied for which it is shown when lack of persistence occurs, i.e. when nor persistent $m$-hubs exist nor persistence occurs.
Before we state our main results, we define the following quantities. Let $(E_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent exponential random variables, where $E_{i}$ has rate $b(i)+d(i)$, and let $(B_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent Bernoulli random variables (also independent of the $E_{i}$), such that
@@
Finally, define
$S_{k}:=\sum_{i=0}^{k-1}E_{i},\qquad\ D:=\inf\{i\in\mathbb{N}_{0}:B_{i}=0\},\qquad\text{and}\qquad\mathcal{R}:=\sum_{j=1}^{D}\delta_{S_{j}},$ (2.6)
@@
$\mu(t):=\lim_{\varepsilon\downarrow 0}\varepsilon^{-1}\mathbb{P}\!\left(\mathcal{R}((t,t+\varepsilon))\geq 1\right),\qquad t\geq 0.$ (2.7)
@@
$\widehat{\mu}(\lambda):=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\sum_{k=1}^{\infty}\prod_{i=0}^{k-1}\frac{b(i)}{b(i)+d(i)+\lambda}.$ (2.8)
The explicit expression for $\widehat{\mu}$ in terms of the sequences $b$ and $d$ follows from [13, Proposition $1.1$]. Then, we define for $k\in\mathbb{N}$ the sequences
$\varphi_{1}(k):=\sum_{i=0}^{k-1}\frac{1}{b(i)+d(i)},\qquad\varphi_{2}(k):=\sum_{i=0}^{k-1}\Big{(}\frac{1}{b(i)+d(i)}\Big{)}^{2},\qquad\rho_{1}(k):=\sum_{i=0}^{k-1}\frac{d(i)}{b(i)+d(i)}.$ (2.9)
@@
In the ‘infinite lifetime’ regime we provide criteria under which persistent $m$-hubs do or do not exist, in the sense of (PH-$m$). In the ‘finite lifetime’ regime we provide criteria under which persistence occurs, in the sense of (P). These results complement the work of the author and Heydenreich in [23] regarding lack of persistence for the PAVD model, in the sense of (LP). The proof techniques for the two regimes are entirely distinct. In the ‘finite lifetime’ regime we use a similar methodology as used in [23], though we now focus on the occurrence of persistence rather than lack of persistence, which requires a finer analysis at times. In the ‘infinite lifetime’ regime we expand and generalise ideas developed by Iyer in [25]. In fact, as we shall see in Section 4, the results in this regime can be proved to hold for more general family of models.
Infinite lifetime regime. We first present our result for the model when vertices can survive forever with positive probability. To this end, we introduce the following assumptions. Recall $\widehat{\mu}$ from (2.8). We assume that
There exists λ>0 such that $\text{There exists $\lambda>0$ such that }\widehat{\mu}(\lambda)<\infty.$ (F-L)
@@
Throughout the paper we assume that Assumption (F-L) “Finite-Laplace” holds. In fact, Assumption (F-L) combined with the continuity of the mapping $\lambda\mapsto\widehat{\mu}(\lambda)$ implies Assumption (N-E) “Non-Explosion”. It guarantees that the tree process is ‘not exploding’, in the sense that there does not exists a unique vertex that obtains an infinite degree, nor a unique path from the root of infinite length. Assumption (C-V) “Converging-Variance” implies that the variance of $S_{k}$ converges with $k$, and is a crucial assumption under which we prove that $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub or that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$. Finally, Assumption (F-D) “Finite-Degree” implies that the random variable $D$, as in (2.6), is finite almost surely (see Lemma 5.1) and governs whether we are in the ‘infinite lifetime’ regime or ‘finite lifetime’ regime. We hence may or may not assume that (F-D) is satisfied.
We then have the following result, where we show under which conditions the tree process $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub or not.
@@
In particular, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent elder in the sense of (PE) ((((i.e. $O<\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely)))) if and only if Assumption (F-D) is not satisfied.
Now, suppose Assumption (N-E) is satisfied and Assumptions (F-D) and (C-V) are not satisfied. Then, for any $m\in\mathbb{N}$, $(T_{n})_{n\in\mathbb{N}}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely. If, instead, Assumption (C-V) is satisfied and, additionally, Assumption (F-L) is satisfied, then for any $m\in\mathbb{N}$ there exists an almost surely finite random variable $I^{(m)}$ such that
@@
That is, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub in the sense of (PH-$m$) $\mathbb{P}_{\mathcal{S}}$-almost surely. Moreover, for any $m\geq 2$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
@@
irrespective of $d$. We additionally need that $d$ is such that Assumption (F-D) is not satisfied to be in the infinite lifetime regime, so that either $d$ converges to zero sufficiently fast or is non-zero at a sufficiently sparse sub-sequence.
@@
Finite lifetime. We now present our result for the PAVD model when vertices die after a finite number of steps almost surely. To this end, we need some additional definitions and assumptions. First, we define the quantity
$R:=\inf_{i\in\mathbb{N}_{0}}(b(i)+d(i)),$ (2.14)
@@
and such that Assumption (F-D) is satisfied. Here, Assumption (F-D) ensures that vertices die after a finite number of steps almost surely. Furthermore, it is crucial that the limit $d^{*}$ is strictly smaller than $R$. Indeed, the ‘finite lifetime’ regime can be divided into two further sub-regimes, called the ‘rich are old’ ‘rich die young’ and regime, based on whether $d^{*}<R$ or $d^{*}>R$ (when assuming that such a $d^{*}$ exists), respectively, see Table 2. The recent work of Heydenreich and the author [23] studies lack of persistence in these regime. They show that, under mild conditions, persistence never occurs in the ‘rich die young’ regime, see [23, Theorem $2.10$]. Our focus here thus lies on the ‘rich are old’ regime, where we show under which conditions persistence does occur.
@@
$\alpha(k):=\rho_{1}(k)-d^{*}\varphi_{1}(k)=\sum_{i=0}^{k-1}\frac{d(i)-d^{*}}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}.$ (2.16)
@@
$\mathcal{K}_{\alpha}(t):=\alpha\big{(}\varphi_{1}^{-1}(t)\big{)}\qquad\text{for }t\geq 0.$ (2.17)
Heuristically, $\mathcal{K}_{\alpha}(t)$ quantifies the difference between $\rho_{1}(\varphi_{1}^{-1}(t))$ and $d^{*}t$, and allows us to precisely quantify probabilities of the form $\mathbb{P}\!\left(D\geq\varphi_{1}^{-1}(t)\right)$ (see Lemma 5.1). We then recall $\widehat{\mu}$ from (2.8). Assumption (F-L) is not sufficient in the ‘finite lifetime’ regime. Rather, we introduce
$\underline{\lambda}:=\inf\{\lambda>0:\widehat{\mu}(\lambda)<\infty\},$ (2.18)

### v2

@@ 1. Introduction
Random graph models based on preferential attachment are quite well-understood nowadays due to a large body of research formed over the last three decades (see [34, 35] and the references therein for a good overview of the literature). These models have become a staple in random graph theory due to their use in modelling real-world networks. Still, a clear gap between most of these models and networks that form in the real world is the fact that preferential attachment graphs allow for growth only. Vertices and edges are sequentially attached to the graph, causing the graph to grow. However, this is not particularly realistic, as nodes and bonds in real-world networks can be removed in almost all contexts as well. As an example, people are born but also pass away; friendships are formed but can also be broken; users of social media can befriend or follow others, but also unfriend or unfollow them and even remove their accounts all together.
@@
Some recent work partially addresses these concerns. Diaz, Lichev, and the author study a model with uniform attachment and uniform vertex removal ($m$ is random, $p_{2}=p_{4}=0$), where the local weak limit, existence of a giant component, and the size and location of the maximum degree are studied [17]. Work of Bellin, Kammerer et al. [5, 6, 26] studies a tree model of uniform attachment and uniform vertex removal ($m=1$, $p_{2}=p_{4}=0$), but where vertices are not removed but are ‘frozen’ and can then no longer make new connections, and where the choice to freeze vertices need not be random. This allows them to study the model in a ‘critical window’ where they obtain precise results for the height of the tree and its scaling limit.
@@
The notion of persistence in evolving random graphs has been studied from different perspectives, starting with the work of Dereich and Mörters [16]. Later, conditions under which persistence and the lack thereof hold have been weakened by Galashin [20], Banerjee and Bhamidi [2], and Iyer [25]. In short (and omitting minor technical assumptions), using the PAVD formulation with $d\equiv 0$, so that there is no vertex death and we recover the classic preferential attachment model, persistence of the maximum degree occurs almost surely if and only if
@@
Heuristically, persistence occurs when high-degree vertices have a sufficiently strong advantage over low-degree vertices when establishing new connections, so that they are most likely to increase their in-degree further. This allows one vertex to attain the largest degree for all but finitely many steps, instead of the largest degree switching between vertices infinitely often. The transition between whether the advantage is strong enough or not lies exactly at the point when the series in (1.3) goes from finite to infinite.
@@
Our contribution. In this article we study the two regimes outlined above. In the infinite lifetime regime, in which vertices can live forever with positive probability, we provide conditions, similar to (1.3), under which a persistent $m$-hub exists almost surely conditionally on the tree process $(T_{n})_{n\in\mathbb{N}}$ surviving, for each $m\in\mathbb{N}$, as well as conditions under which persistent hubs do not exist. This analysis generalises and expands recent work of Iyer [25], which focusses on the case $d\equiv 0$ and $m=1$.
@@
The second regime is the finite lifetime regime, in which vertices are killed after a finite number of steps almost surely. Here, we provide conditions, similar to (1.3), under which persistence occurs. This builds on recent work by the author and Heydenreich [23], in which lack of persistence is studied for the PAVD model in both the infinite and the finite lifetime regime, and extends the methodology developed by Deijfen in [13] using an embedding of the discrete tree process in a continuous-time branching process known as a Crump-Mode-Jagers branching process. We use a more precise formulation of this embedding compared to Deijfen, which allows us to simplify certain proofs of results in [13], but mainly to develop novel and precise results regarding the behaviour of the branching process embedding that were unattainable previously. Furthermore, we extend ideas used by Banerjee and Bhamidi in [2] to precisely analyse the optimal window in which the vertex with the largest degrees is found in the finite lifetime regime.
@@
Structure of the paper. We state the main results and the necessary assumptions in Section 2. The methodology used in the analysis in presented in Section 3. Section 4 studies the existence of persistent $m$-hubs in the infinite lifetime regime, whereas Section 5 focusses on the occurrence of persistence in the finite lifetime regime. We conclude with a discussion of our work, state several open problems, and discuss possibilities for future directions in Section 6.
@@
Notation. Throughout the paper we use the following notation: we let $\mathbb{N}\coloneq\{1,2,\ldots\}$ denote the natural numbers, set $\mathbb{N}_{0}\coloneq\{0,1,\ldots\}$ to include zero and let $t]\coloneq\{i\in\mathbb{N}:i\leq t\}$ for any $t\geq 1$. For $x,y\in\mathbb{R}$, we let $\lceil x\rceil\coloneq\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor\coloneq\sup\{n\in\mathbb{Z}:n\leq x\}$, and let $x\wedge y\coloneq\min\{x,y\}$ and $x\vee y\coloneq\max\{x,y\}$. For sequences $(a_{n})_{n\in\mathbb{N}},(b_{n})_{n\in\mathbb{N}}$ such that $b_{n}$ is positive for all $n$ we say that $a_{n}=o(b_{n})$ and $a_{n}=\mathcal{O}(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$ and if there exists a constant $C>0$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$, respectively. We write $a_{n}=\Theta(b_{n})$ if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$. For random variables $X,(X_{n})_{n\in\mathbb{N}}$, and $Y$ we let $X_{n}\overset{\mathrm{d}}{\longrightarrow}X,X_{n}\overset{\mathbb{P}}{\longrightarrow}X$ and $X_{n}\overset{\mathrm{a.s.}}{\longrightarrow}X$ denote convergence in distribution, probability and almost sure convergence of $X_{n}$ to $X$, respectively. $X\preceq Y$ (resp. $X\succeq Y$) denotes that $X$ is stochastically dominated by $Y$ (resp. $X$ stochastically dominates $Y$). For a sequence $(\mathcal{E}_{t})_{t\in\mathbb{I}}$ of events, where $\mathbb{I}=\mathbb{N}$ or $\mathbb{I}=[0,\infty)$, we say that $\mathcal{E}_{t}$ holds with high probability when $\lim_{t\to\infty}\mathbb{P}\!\left(\mathcal{E}_{t}\right)=1$. Finally, for an event $\mathcal{S}$ such that $\mathbb{P}\!\left(\mathcal{S}\right)>0$ we let $\mathbb{P}_{\mathcal{S}}\left(\cdot\right)\coloneq\mathbb{P}(\cdot\,|\,\mathcal{S})$ and $\mathbb{E}_{\mathcal{S}}{}{\cdot}\coloneq\mathbb{E}\left[\cdot\,|\,\mathcal{S}\right]$ denote the conditional probability measure and conditional expectation, respectively. Then, we let $\xrightarrow{\mathbb{P}_{\mathcal{S}}-\mathrm{a.s.}}$ respectively $\overset{\mathbb{P}_{\mathcal{S}}}{\longrightarrow}$ denote almost sure convergence and convergence in probability with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$.
@@ 2. Results
$\mathcal{S}\coloneq\bigcap_{n\in\mathbb{N}}\{\mathcal{A}_{n}\neq\varnothing\}$ (2.1)
@@
$I_{n}^{(1)}\coloneq\min\{i\in\mathcal{A}_{n}\colon\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\},$ (2.3)
@@
$I_{n}^{(m)}\coloneq\min\{i\in\mathcal{A}_{n}\setminus\{I_{n}^{(i)}\}_{i\in[m-1]}\colon\deg_{n}(i)\geq\deg_{n}(j)\text{ for all }j\in\mathcal{A}_{n}\setminus\{I_{n}^{(i)}\}_{i\in[m-1]}\}.$ (2.4)
@@
We are interested in the long-term behaviour of $I_{n}^{(m)}$ and $O_{n}$. Since $O_{n}$ is non-decreasing in $n$, it follows that $O_{n}$ converges $\mathbb{P}_{\mathcal{S}}$-almost surely to a limit $O$. We say that a persistent elder exists when
$O\coloneq\lim_{n\to\infty}O_{n}\text{ is finite }\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (PE)
When $O<\infty$ holds, we can think of $O$ as the first vertex that is added to the tree that will never be killed. It is clear that this limit exists if and only if vertices can live forever with positive probability (i.e. by never being both selected and killed). Otherwise, $O=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. In a similar manner, we say that a persistent $m$-hub exists when there exists a $\mathbb{P}_{\mathcal{S}}$-almost surely finite random variable $I^{(m)}$ such that
$\lim_{n\to\infty}I_{n}^{(m)}=I^{(m)}\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (PH-$m$)
Intuitively, the existence of the limit in (PH-$m$) implies that the vertex with the $m^{\mathrm{th}}$ largest degree eventually stabilizes, in the sense that vertex $I^{(m)}$ attains the $m^{\mathrm{th}}$ largest degree in $T_{n}$ for all but finitely many $n$.
By noting that that $I_{n}^{(m)}\geq O_{n}$ for all $m,n\in\mathbb{N}$, it follows that the limit $I^{(m)}$ in (PH-$m$) cannot exist when the limit $O$ in (PE) equals infinity. That is, if no persistent elder exists, then also no persistent $m$-hub exists. In this case, we instead focus on the growth rate of the variables $I_{n}^{(m)}$ and $O_{n}$. We say that persistence occurs in $(T_{n})_{n\in\mathbb{N}}$ when
@@
We intuitively think of (P) as the vertex with the largest degree being ‘close’ to the oldest alive vertex, in terms of their label, whereas (LP) implies that they are ‘far apart‘.
Clearly, when a persistent $1$-hub exists, as in (PH-$m$) with $m=1$, then the existence of a persistent elder, as in (PE), and the existence of a persistent $1$-hub, as in (P), are implied, since $1\leq O_{n}\leq I_{n}^{(1)}$ and $O_{n}$ is non-decreasing. Similarly, lack of persistence, as in (LP), implies that no persistent $1$-hub exists.
@@
  • A class of PAVD models for which we prove that it does not contain a persistent $m$-hub for any $m\in\mathbb{N}$, but persistence does occur. (Theorem 2.6)
To be able to state our main results, we need to define the following quantities. Let $(E_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent exponential random variables, where $E_{i}$ has rate $b(i)+d(i)$, and let $(B_{i})_{i\in\mathbb{N}_{0}}$ be a sequence of independent Bernoulli random variables (also independent of the $E_{i}$), such that
@@
Then, define
$S_{k}\coloneq\sum_{i=0}^{k-1}E_{i},\qquad\ D\coloneq\inf\{i\in\mathbb{N}_{0}:B_{i}=0\},\qquad\text{and}\qquad\mathcal{R}\coloneq\sum_{j=1}^{D}\delta_{S_{j}},$ (2.6)
@@
$\mu(t)\coloneq\lim_{\varepsilon\downarrow 0}\varepsilon^{-1}\mathbb{P}\!\left(\mathcal{R}((t,t+\varepsilon))\geq 1\right),\qquad t\geq 0.$ (2.7)
@@
$\widehat{\mu}(\lambda)\coloneq\int_{0}^{\infty}\mathrm{e}^{-\lambda t}\mu(t)\,\mathrm{d}t=\mathbb{E}\left[\sum_{k=1}^{D}\exp\bigg(-\lambda\sum_{i=0}^{k-1}E_{i}\bigg)\right]=\sum_{k=1}^{\infty}\prod_{i=0}^{k-1}\frac{b(i)}{b(i)+d(i)+\lambda}.$ (2.8)
The explicit expression for $\widehat{\mu}$ in terms of the sequences $b$ and $d$ follows from [13, Proposition $1.1$]. Finally, we define for $k\in\mathbb{N}$ the sequences
$\varphi_{1}(k)\coloneq\sum_{i=0}^{k-1}\frac{1}{b(i)+d(i)},\quad\varphi_{2}(k)\coloneq\sum_{i=0}^{k-1}\Big(\frac{1}{b(i)+d(i)}\Big)^{2},\quad\rho_{1}(k)\coloneq\sum_{i=0}^{k-1}\frac{d(i)}{b(i)+d(i)}.$ (2.9)
@@
The proof techniques for the results related to each regime are entirely distinct. In the ‘finite lifetime’ regime we use a similar methodology as used in [23], though we now focus on the occurrence of persistence rather than lack of persistence, which requires a finer analysis at times. In the ‘infinite lifetime’ regime we expand and generalise ideas developed by Iyer in [25].
Infinite lifetime regime. We first present our result for the PAVD model when vertices can survive forever with positive probability. To this end, we introduce the following assumptions. Recall $\widehat{\mu}$ from (2.8). We assume that
$\text{There exists $\lambda>0$ such that }\widehat{\mu}(\lambda)<\infty.$ (F-L)
@@
Throughout the paper we assume that Assumption (F-L) “Finite-Laplace” holds. It ensures that the branching process used in the analysis does not grow infinitely large in finite time. Assumption (N-E) “Non-Explosion” is of a similar nature and implies that there does not exist a unique vertex in the tree that obtains an infinite degree (as $n\to\infty$). Assumption (C-V) “Converging-Variance” implies that the variance of $S_{k}$ converges with $k$, and is a crucial assumption under which we prove that a persistent $m$-hub exists (as in (PH-$m$)) or that persistence occurs (as in (P)). Finally, Assumption (F-D) “Finite-Degree” implies that the random variable $D$, as in (2.6), is finite almost surely (see Lemma 5.1) and governs whether we are in the ‘infinite lifetime’ regime or ‘finite lifetime’ regime. We hence may or may not assume that (F-D) is satisfied.
Deijfen [13] shows that Assumption (F-L) implies that either Assumption (N-E) or Assumption (F-D) must be satisfied. We do not include an analysis of the case in which Assumption (N-E) is not satisfied but Assumption (F-D) is, which is on-going work.
We then have the following result, where we show under which conditions a persistent $m$-hub does or does not exist.
@@
In particular, a persistent elder exists, in the sense of (PE), if and only if Assumption (F-D) is not satisfied.
Suppose Assumption (N-E) is satisfied and Assumptions (F-D) and (C-V) are not satisfied. Then, for any $m\in\mathbb{N}$, a persistent $m$-hub does not exist $\mathbb{P}_{\mathcal{S}}$-almost surely.
If Assumptions (N-E), (C-V), and (F-L) are satisfied and Assumption (F-D) is not, then for any $m\in\mathbb{N}$ there exists an almost surely finite random variable $I^{(m)}$ such that
@@
That is, a persistent $m$-hub exists $\mathbb{P}_{\mathcal{S}}$-almost surely, in the sense of (PH-$m$). Moreover, for any $m\geq 2$ there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
@@
irrespective of $d$. We additionally need that $d$ is such that Assumption (F-D) is not satisfied to be in the infinite lifetime regime, so that either $d$ converges to zero sufficiently fast or is non-zero only at a sufficiently sparse sub-sequence.
@@
Finite lifetime. We now present our result for the PAVD model when vertices die after a finite number of steps almost surely. The results in Theorem 2.1 are qualitative, as they state whether certain random variables converge or diverge. Our result in finite lifetime, on the other hand, is quantitative. It shows that certain random variables grow at the same rate (in the sense of (P)). To this end, we need some additional definitions and assumptions to more finely control the behaviour of the model. First, we define
$R\coloneq\inf_{i\in\mathbb{N}_{0}}(b(i)+d(i)),$ (2.14)
@@
and such that Assumption (F-D) is satisfied. Here, Assumption (F-D) ensures that vertices die after a finite number of steps almost surely. Furthermore, it is crucial that the limit $d^{*}$ is strictly smaller than $R$. Indeed, the ‘finite lifetime’ regime can be divided into two further sub-regimes, called the ‘rich are old’ ‘rich die young’ and regime, based on whether $d^{*}<R$ or $d^{*}>R$ (when assuming that such a $d^{*}$ exists), respectively, see Table 2. The recent work of Heydenreich and the author [23] studies lack of persistence in these regime. They show that, under mild conditions, persistence never occurs in the ‘rich die young’ regime, see [23, Theorem $2.10$], and that lack of persistence occurs in the ‘rich are old’ regime when Assumption (C-V) is not satisfied. Our focus here thus lies on the ‘rich are old’ regime, where we show that persistence does occur when Assumption (C-V) is satisfied (and additional technical assumptions).
@@
$\alpha(k)\coloneq\rho_{1}(k)-d^{*}\varphi_{1}(k)=\sum_{i=0}^{k-1}\frac{d(i)-d^{*}}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}.$ (2.16)
@@
$\mathcal{K}_{\alpha}(t)\coloneq\alpha\big(\varphi_{1}^{-1}(t)\big)\qquad\text{for }t\geq 0.$ (2.17)
Heuristically, $\mathcal{K}_{\alpha}(t)$ quantifies the difference between $\rho_{1}(\varphi_{1}^{-1}(t))$ and $d^{*}t$, and allows us to precisely quantify probabilities of the form $\mathbb{P}\!\left(D\geq\varphi_{1}^{-1}(t)\right)$ (see Lemma 5.1). We then recall $\widehat{\mu}$ from (2.8). Assumption (F-L) is not sufficiently strong for the proof techniques used in the ‘finite lifetime’ regime. Rather, we introduce
$\underline{\lambda}\coloneq\inf\{\lambda>0:\widehat{\mu}(\lambda)<\infty\},$ (2.18)